OpenStudy (anonymous):

A few of my answers aren't matching up with my answer key. I was wondering if I'm just making careless errors or if the answer key may be wrong.

OpenStudy (anonymous):

\[-14-3(2x+1)=8x+1\]\[-14-6x-3=8x+1\]\[-17-6x=8x+1\]\[-17=14x+1\]\[-18=14x\]\[x=\frac{ -18 }{ 14 }=\frac{ -9 }{ 7 }\]

Nnesha (nnesha):

your answer is correct

OpenStudy (anonymous):

Okay thanks ^^

Nnesha (nnesha):

substitute x for -1 \[-14-3(2(-1)+1)=8(-1)+1\] -14-3(-1)=-8+1 -14+3=-7 -11=-7 both sides aren't equal so -1 is wrong